OpenStudy (anonymous):

The polynomial given below has _____ root(s). 2x2 + 5x + 2 A. two complex B. two positive C. two negative D. one positive and one negative

OpenStudy (amistre64):

we can start with the determinant to determine the nature of the roots

OpenStudy (amistre64):

b^2 - 4ac 25 -4(2)(2) = 25 -16 = 9

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

oh no C.

OpenStudy (amistre64):

it has at 2 real roots at least

OpenStudy (anonymous):

x=-2,-1/2

OpenStudy (amistre64):

yep, a graph confirms at least they are 2 negatives ;) http://www.wolframalpha.com/input/?i=2x2+%2B+5x+%2B+2